OpenStudy (wild_child117):

Solve for x. −7 ≥ 13 − 5x _____________________ a. x ≥ −4 b. x ≤ 4 c. x ≥ 4 d. x ≤ −4

OpenStudy (wild_child117):

@Will.H

OpenStudy (will.h):

\[-7 - 13 \ge - 5x\] \[-20 \ge -5x\] \[4 \le x\]

OpenStudy (will.h):

so C again

OpenStudy (wild_child117):

thank you i have two more is that okay in separate post?

OpenStudy (will.h):

sure